# Optimizing for ROC Curves on Class-Imbalanced Data by Training over a Family of Loss Functions

### v1

@@ Abstract
Although binary classification is a well-studied problem in computer vision, training reliable classifiers under severe class imbalance remains a challenging problem. Recent work has proposed techniques that mitigate the effects of training under imbalance by modifying the loss functions or optimization methods. While this work has led to significant improvements in the overall accuracy in the multi-class case, we observe that slight changes in hyperparameter values of these methods can result in highly variable performance in terms of Receiver Operating Characteristic (ROC) curves on binary problems with severe imbalance. To reduce the sensitivity to hyperparameter choices and train more general models, we propose training over a family of loss functions, instead of a single loss function. We develop a method for applying Loss Conditional Training (LCT) to an imbalanced classification problem. Extensive experiment results, on both CIFAR and Kaggle competition datasets, show that our method improves model performance and is more robust to hyperparameter choices. Code will be made available here: https://github.com/klieberman/roc_lct.
Class Imbalance, ROC Curves, Machine Learning
@@ 1 Introduction
Figure 1: Distribution of Area Under the ROC Curve (AUC) values obtained by training the same model on the SIIM-ISIC Melanoma classification dataset with 48 different combinations of hyperparameters on VS loss. Results are shown at three different imbalance ratios. As the imbalance becomes more severe, model performance drops and the variance in performance drastically increases. LCT addresses both of these issues by training over a family of loss functions, instead of a single loss function.
Consider a classifier that takes images of skin lesions and predicts whether the lesions are melanoma (Rotemberg et al., 2020). Such a system could be especially valuable in underdeveloped countries where expert resources for diagnosis are scarce (Cassidy et al., 2022). Classifying melanoma from images is a problem with class imbalance since benign lesions are far more common than melanomas. Furthermore, the accuracy on the melanoma (minority) class is much more important than the accuracy on the benign (majority) class because predicting a benign lesion as melanoma would result in the cost of a biopsy while predicting a melanoma lesion as benign could result in the melanoma spreading before the patient can receive appropriate treatment.
In this case, overall accuracy, even on a balanced test set, is clearly an inadequate metric, as it implies that the accuracies of both classes are equally important. Instead, Receiver Operating Characteristic (ROC) curves are better suited for such problems (Egan, 1975). These curves plot the tradeoff between the true positive rate (TPR) on the y-axis and the false positive rate (FPR) on the x-axis over a range of classification thresholds. Unlike scalar metrics (e.g., overall accuracy on a balanced test set or $F_{\beta}$), ROC curves show model performance over a wide range of classification thresholds. This allows practitioners to understand how the model’s performance changes based on different classification thresholds and choose the best tradeoff for their needs. ROC curves can also be summarized by their Area Under the Curve (AUC) (Egan, 1975). Furthermore, both ROC curves and AUC have several mathematical properties that make them preferred to alternative precision-recall curves (Flach & Kull, 2015) [^1] We provide definitions and visualizations of commonly-used metrics for binary problems with imbalanced data in Appendix A.
@@
In particular, we adapt models trained with Vector Scaling (VS) loss. This loss is a general loss function that encompasses many other loss functions for imbalanced data as special cases (Kini et al., 2021). It modifies Cross-entropy loss by introducing additive and multiplicative factors to the logits, which changes the inter-class margins. Although VS loss has shown strong performance in the multi-class setting, it does have hyperparameters that require tuning (e.g., the additive and multiplicative factors on the loss function).
We find that, in the binary case, the effect of these hyperparameters is small and reasonable at moderate imbalance ratios $\beta$ (where $\beta$ is defined as the ratio of majority to minority samples); however, at large imbalance ratios, small differences in these hyperparameters lead to very wide variance in the results (Figure 1). This figure shows that increasing the imbalance ratio not only decreases the AUCs (as expected) but also drastically increases the variance in AUCs obtained by training models with slightly different hyperparameter values. This suggests that traditional models trained using losses with a fixed hyperparameter set are especially noisy, and extensive hyperparameter tuning is necessary in the case of large imbalance ratios.
To address this problem and train a more robust model, we propose training a single model on a wide range of hyperparameter values using Loss Conditional Training (Dosovitskiy & Djolonga, 2020). We design a method for applying LCT to classification problems under imbalance via VS loss. We find that this method not only reduces the variance in model performance caused by hyperparameter choices but also improves performance over the best hyperparameter choices since it optimizes for many tradeoffs on the ROC curve. We provide extensive results, both on CIFAR datasets and datasets of real applications derived from Kaggle competitions, at multiple imbalance ratios and across a wide range of hyperparameter choices.
@@
  • We identify that higher levels of imbalance are not only associated with worse model performance but also more variance.
  • We propose training over a family of loss functions as a way to reduce this variance.
  • We modify Loss Conditional Training (LCT) for binary classification under imbalance.
  • We show that this method consistently improves performance and reduces variance at high imbalance ratios.
@@ 2 Related work
Many solutions have been proposed to address class imbalance, including several specialized loss functions and optimization methods (Cao et al., 2019; Rangwani et al., 2022; Buda et al., 2018; Kini et al., 2021; Shwartz-Ziv et al., 2023). Perhaps the simplest of these is to change the class weights in the loss function so that the minority and majority classes have “balanced” class weights or class weights that are inversely proportional to the frequency of the class in the training set (e.g., weighted Cross-entropy loss (Xie & Manski, 1989)). Another popular loss function is Focal loss, which down-weights “easy” samples (i.e., samples with high predictive confidence) (Lin et al., 2017).
More recently, several loss functions have been proposed that add additive and multiplicative factors to the logits before they are input to the softmax function (Cao et al., 2019; Menon et al., 2021; Ye et al., 2020). These aim to enforce larger margins on the minority class and/or calibrate the models. Kini et al. (2021) recognized that many of the previous loss functions for addressing class imbalance can be expressed by one general loss function: Vector Scaling (VS) loss which gives strong performance on multi-class datasets after hyperparameter tuning (Kini et al., 2021). Du et al. (2023) use a global and local mixture consistency loss, contrastive learning, and a dual head architecture. To gain insights into the effects of training over a family of hyperparameters, we apply our method to standard networks and VS loss, which has hyperparameters with well-known effects.
Additionally, Rangwani et al. (2022) proposed using Sharpness Aware Minimization (SAM)(Foret et al., 2021) as an alternative optimizer and found that this helped the model escape saddle points in multi-class problems with imbalance. Similarly, Shwartz-Ziv et al. (2023) identify several tweaks that can be made to models—including batch size, data augmentation, specialized optimizers, and label smoothing—which can all improve training.
These methods are predominantly tested on multi-class datasets with overall accuracy on a balanced test set as the primary metric. We instead propose to optimize for the ROC curve in the binary case by training over a family of loss functions. To this end, we use Loss Conditional Training (LCT), which was proposed as a way to train one model to work over several rates in a variable-rate problem (Dosovitskiy & Djolonga, 2020). LCT was proposed for applications such as neural image compression and variational autoencoders and, to our knowledge, we are the first to use it to improve ROC curves of a classification problem. [^2] We provide a more detailed related works section in  Appendix B
@@ 3.2 Predictor and predictions
Let $f$ be a predictor with weights $\boldsymbol{\theta}$ and $\mathbf{z}=f(\mathbf{x},\boldsymbol{\theta})$ be $f$’s output for input $\mathbf{x}$. We assume $\mathbf{z}$ are logits (i.e., a vector $(z_{0},z_{1})$ of unnormalized scalars where $z_{c}$ is the logit corresponding to class $c$) and $\mathbf{p}$ are the outputs of the softmax function (i.e., a vector $(p_{0},p_{1})$ of normalized scalars such that $p_{0}+p_{1}=1$). Then the model’s prediction is
@@ 3.3 Loss function
For all experiments, we use the Vector Scaling (VS) loss as defined by Kini et al. (2021). This loss is a modification of weighted cross entropy loss with two hyperparameters $\Delta,\iota$ that specify an affine transformation $\Delta z+\iota$ for each logit:
@@
$\iota_{c}=\tau\log\left(\frac{n_{c}}{n}\right)\quad\text{and}\quad\Delta_{c}=\left(\frac{n_{c}}{n_{0}}\right)^{\gamma}.$ (3)
@@
where $\Omega\in[0,1]$. In the binary case, we can simplify the loss as follows (see Appendix G for details):
@@
Note that VS loss is equivalent to Cross-entropy loss when $\Omega=0.5,\gamma=0$, and $\tau=0$. It is also equal to weighted cross entropy loss when $\Omega=\frac{n_{0}}{n},\gamma=0$, and $\tau=0$.
@@ 4 Analysis on the effects of hyperparameters
Figure 2: Distribution of results obtained from training 512 models with different hyperparameter values. Left: mean, min, max, and 95% confidence interval of ROC curves. Right: boxplot of distribution of Area Under the ROC Curve (AUC) values. Under class imbalance, differences in hyperparameter values can lead to a large variance in results.
Refer to caption Metric $R^{2}$ AUC 0.30 Acc. 0.61 TPR 0.92
@@ 4.1 Variance of results
As discussed in  Section 3.3, our parameterization of VS loss has three hyperparameters: $\Omega,\gamma$ and $\tau$. Although there are intuitive and theoretical rationale for each of these hyperparameters (Kini et al., 2021), it is not clear how sensitive a model’s performance is to their values, especially in terms of ROC curves on a binary dataset.
@@ 4.2 Correlation between hyperparameter values and scalar metrics
To test this, we fit polynomials between the three hyperparameters ($\Omega,\gamma,\tau$) and two scalar metrics: AUC and overall accuracy on a balanced test set when $t=0.5$. Although TPR, equivalently recall, is not a good indicator of a model’s overall performance by itself (because it is trivial to increase TPR at the expense of FPR), we also fit a polynomial to TPR to gain insights on the effect of the hyperparameters on a ROC curve. We show results, including the $R^{2}$ of these regressions, in the three plots in Figure 3. We also fit three (one for AUC, accuracy, and TPR) degree-2 polynomials using all three hyperparameters as features and report $R^{2}$ values of these polynomials in the table in Figure 3. We choose polynomials of degree two instead of linear regression because we expect that an optimum exists.
We see that no single hyperparameter is strongly correlated with AUC; however, 30% of the variance can be explained by the polynomial fit to all three hyperparameters. This suggests that some variance can be removed by choosing strong hyperparameters; however, much of the variance comes from other sources. Additionally, only 61% of the variance in overall accuracy can be explained by the polynomial with all three hyperparameters. While this is significantly better than AUC’s $R^{2}$, it is still far from perfect. Thus, hyperparameter values do not affect performance metrics in a very predictable way. It would be preferable to reduce the variance of VS Loss over performing extensive hyperparameter tuning.
@@ 4.3 Connection between hyperparameter values and softmax outputs
The third plot in Figure 3 shows that $\tau$ is strongly correlated with TPR when the threshold $t=0.5$ ($R^{2}=0.92$). To understand this, recall that $\tau$ affects the additive term on the logits in Equations 2 and 3. Specifically, increasing $\tau$ enforces a larger margin on the minority class. In general, this will lead to increasing the softmax scores of the minority class at inference time, which is similar to post-hoc VS calibration (Zhao et al., 2020). Thus, when $t$ is constant, increasing $\tau$ will likely cause an increase in TPR.
We can also visualize the effect of $\tau$ by considering the set of “break-even” logits, which are defined as the set of points of $\mathbf{z}$ where the loss is equal whether the sample has label 0 or 1 (i.e., the points of $\mathbf{z}$ such that $\ell_{VS}(1,\mathbf{z})=\ell_{VS}(0,\mathbf{z})$). With regular Cross-entropy loss, this set is $z_{1}=z_{0}$. Assume instead that we keep $\Omega=0.5,\gamma=0$, but have arbitrary values for $\tau$, then the equation for the set of break-even points becomes $z_{1}=z_{0}+\tau\log\beta$. In other words, varying $\tau$ is equivalent to shifting the set of break-even points (and the rest of the loss landscape) by $\tau\log(\beta)$. Figure 4 visualizes this. [^4] Varying $\tau$ also shifts the loss landscape in a similar way for any fixed $\Omega\in(0,1)$ and $\gamma>0$. See Appendix I for details.
In addition, $\Omega$ and $\gamma$ also have well-understood effects on the logit outputs.  Appendices I and J include a general equation for the set of break-even points and more loss contours. Just as we saw in the example with $\tau$, different combinations of hyperparameters have different effects on how $\mathbf{z}$ is scaled, which in turn affects their softmax scores. Thus, training with different values of hyperparameters corresponds to optimizing for different tradeoffs of TPR and FPR.
Figure 4: Effect of $\tau$ on loss landscape. Each plot shows $\ell_{VS}(1,\mathbf{z})-\ell_{VS}(0,\mathbf{z})$ over $z_{0},z_{1}\in[-5,5]$ for $\beta=10$. White: “break-even” points. $\boldsymbol{\tau}$ shifts the loss landscape.

### v2

@@ Abstract
Although binary classification is a well-studied problem in computer vision, training reliable classifiers under severe class imbalance remains a challenging problem. Recent work has proposed techniques that mitigate the effects of training under imbalance by modifying the loss functions or optimization methods. While this work has led to significant improvements in the overall accuracy in the multi-class case, we observe that slight changes in hyperparameter values of these methods can result in highly variable performance in terms of Receiver Operating Characteristic (ROC) curves on binary problems with severe imbalance. To reduce the sensitivity to hyperparameter choices and train more general models, we propose training over a family of loss functions, instead of a single loss function. We develop a method for applying Loss Conditional Training (LCT) to an imbalanced classification problem. Extensive experiment results, on both CIFAR and Kaggle competition datasets, show that our method improves model performance and is more robust to hyperparameter choices. Code is available at https://github.com/klieberman/roc_lct.
@@ 1 Introduction
Consider a classifier which takes images of skin lesions and predicts whether the lesions are melanoma Rotemberg et al. (2020). Such a system could be especially valuable in underdeveloped countries where expert resources for diagnosis are scarce Cassidy et al. (2022). Classifying melanoma from images is a problem with class imbalance since benign lesions are far more common than melanomas. Furthermore, the accuracy on the melanoma (minority) class is much more important than the accuracy on the benign (majority) class because predicting a benign lesion as melanoma would result in the cost of a biopsy while predicting a melanoma lesion as benign could result in the melanoma spreading before the patient can receive appropriate treatment.
In this case, overall accuracy, even on a balanced test set, is clearly an inadequate metric, as it implies that the accuracies on both classes are equally important. Instead, Receiver Operating Characteristic (ROC) curves are better suited for such problems Egan (1975). These curves plot the tradeoff between the true positive rate (TPR) on the y-axis and the false positive rate (FPR) on the x-axis over a range of classification thresholds. Unlike scalar metrics (e.g., overall accuracy on a balanced test set or $F_{\beta}$), ROC curves show model performance over a wide range of classification thresholds. This allows practitioners to understand how the model’s performance changes based on different classification thresholds and choose the best tradeoff for their needs. ROC curves can also be summarized by their Area Under the Curve (AUC) Egan (1975). Furthermore, both ROC curves and AUC have several mathematical properties which make them preferred to alternative precision-recall curves Flach & Kull (2015) [^1] We provide definitions and visualizations of commonly-used metrics for binary problems with imbalanced data in Appendix A.
@@
Figure 1: Distribution of Area Under the ROC Curve (AUC) values obtained by training the same model on the SIIM-ISIC Melanoma classification dataset with 48 different combinations of hyperparameters on VS Loss (hyperparameter values are given in Section 6). Results are shown at three different imbalance ratios. As the imbalance becomes more severe, model performance drops and the variance in performance drastically increases. LCT addresses both of these issues by training over a family of loss functions, instead of a single loss function with one combination of hyperparameter values.
In particular, we adapt Vector Scaling (VS) loss, which is a general loss function for imbalanced learning with strong theoretical backing Kini et al. (2021). VS loss is a modification of Cross-entropy loss that adjusts the logits via additive and multiplicative factors. There is theory supporting the use of both of these factors: multiplicative factors are essential for the terminal phase of training, but these have negative effects early during training, so additive factors are necessary speed up convergence. Although VS loss has shown strong performance in the multi-class setting, it does have hyperparameters which require tuning (e.g., the additive and multiplicative factors on the loss function).
We find that, in the binary case, the effect of these hyperparameters is small and reasonable at moderate imbalance ratios $\beta$ (where $\beta$ is defined as the ratio of majority to minority samples); however, at large imbalance ratios, small differences in these hyperparameters lead to very wide variance in the results (Figure 1). This figure shows that increasing the imbalance ratio not only decreases the AUCs (as expected), but also drastically increases the variance in AUCs obtained by training models with slightly different hyperparameter values.
In this work, we highlight the practical effect of the theoretically-motivated VS loss on the ROC metric, especially for data problems with high imbalance ratios. We propose a method that adapts VS loss to align the training objective more closely with ROC curve optimization. Our method trains a single model on a wide range of hyperparameter values using Loss Conditional Training Dosovitskiy & Djolonga (2020). We find that this method not only reduces the variance in model performance caused by hyperparameter choices, but also improves performance over the best hyperparameter choices since it optimizes for many tradeoffs on the ROC curve. We provide extensive results– both on CIFAR datasets and datasets of real applications derived from Kaggle competitions– at multiple imbalance ratios and across a wide range of hyperparameter choices.
@@
  • We identify that higher levels of imbalance are not only associated with worse model performance, but also more variance.
  • We recognize that training over a range of hyperparameter values can actually benefit classification models that are otherwise prone to overfitting to a single loss function.
  • We propose using Loss Conditional Training (LCT) to improve the training regimen for classification models trained under imbalance.
  • We show that this method consistently improves performance at high imbalance ratios.
@@ 2 Related work
Many solutions have been proposed to address class imbalance, including several specialized loss functions and optimization methods Cao et al. (2019); Rangwani et al. (2022); Buda et al. (2018); Kini et al. (2021); Shwartz-Ziv et al. (2023). Perhaps the simplest of these is to change the class weights in the loss function so that the minority and majority class have “balanced” class weights or class weights which are inversely proportional to the frequency of the class in the training set (e.g., weighted cross-entropy loss Xie & Manski (1989)). Another popular loss function is Focal loss, which down-weights “easy” samples (i.e., samples with high predictive confidence) Lin et al. (2017).
More recently, several loss functions have been proposed which add additive and multiplicative factors to the logits before they are input to the softmax function Cao et al. (2019); Ye et al. (2020); Menon et al. (2021). These aim to enforce larger margins on the minority class and/or calibrate the models. Kini et al. (2021) recognized that many of the previous loss functions for addressing class imbalance can be expressed by one general loss function: Vector Scaling (VS) loss, which gives strong performance on multi-class datasets after hyperparameter tuning Kini et al. (2021). Du et al. (2023) use a global and local mixture consistency loss, contrastive learning, and a dual head architecture.
Additionally, Rangwani et al. (2022) proposed using Sharpness Aware Minimization (SAM) as an alternative optimizer and found that this helped the model escape saddle points in multi-class problems with imbalance Foret et al. (2021). Similarly, Shwartz-Ziv et al. (2023) identify several tweaks that can be made to models—including batch size, data augmentation, specialized optimizers, and label smoothing—which can all improve training.
These methods are predominantly tested on multi-class datasets with overall accuracy on a balanced test set as the primary metric. We instead propose to optimize for the ROC curve in the binary case by training over a family of loss functions. To this end, we use Loss Conditional Training (LCT), which was proposed as a way to train one model to work over several rates in a variable-rate problem Dosovitskiy & Djolonga (2020). LCT was proposed for applications such as neural image compression and variational autoencoders and, to our knowledge, we are the first to use it to improve ROC curves of a classification problem.
@@ 3.2 Predictor and predictions
Let $f$ be a predictor with weights $\bm{\theta}$ and $\mathbf{z}=f(\mathbf{x},\bm{\theta})$ be $f$’s output for input $\mathbf{x}$. We assume $\mathbf{z}$ are logits (i.e., a vector $(z_{0},z_{1})$ of unnormalized scalars where $z_{c}$ is the logit corresponding to class $c$) and $\mathbf{p}$ are the outputs of the softmax function (i.e., a vector $(p_{0},p_{1})$ of normalized scalars such that $p_{0}+p_{1}=1$). Then the model’s prediction is
@@ 3.3 Loss function
For all experiments, we use the Vector Scaling (VS) loss as defined by Kini et al. (2021). This loss is a modification of weighted Cross-entropy loss with two hyperparameters $\Delta,\iota$ that specify an affine transformation $\Delta z+\iota$ for each logit:
@@
$\iota_{c}=\tau\log\left(\frac{n_{c}}{n}\right)\quad\text{and}\quad\Delta_{c}=\left(\frac{n_{c}}{n_{0}}\right)^{\gamma}$ (3)
@@
where $\Omega\in[0,1]$. In the binary case, we can simplify the loss as follows (see Appendix E for details):
@@
Note that VS loss is equivalent to cross entropy loss when $\Omega=0.5,\gamma=0$, and $\tau=0$. It is also equal to weighted cross entropy loss when $\Omega=\frac{n_{0}}{n},\gamma=0$, and $\tau=0$.
@@ 4 Analysis on the effects of hyperparameters
Figure 2: Distribution of results obtained from training 512 models with different hyperparameter values. Left: mean, min, max, and 95% confidence interval of ROC curves. Right: boxplot of distribution of Area Under the ROC Curve (AUC) values.
Metric $R^{2}$ AUC 0.11 Acc. 0.60 TPR 0.95
@@ 4.1 Variance of results
As discussed in Section 3.3, our parameterization of VS loss has three hyperparameters: $\Omega,\gamma$ and $\tau$. Each of these has theory associated with its effect on training under imbalance. $\Omega$ controls the overall balance between the magnitude of the gradients from the minority and majority class. $\gamma$ was originally proposed to compensate for the difference between the magnitude of the minority-class logits at training and testing time Ye et al. (2020); however, Kini et al. (2021) showed that it is essential for training beyond zero error and optimizing with $\gamma>0$ is equivalent to a solving Cost-Sensitive Support Vector Machine problem. $\tau$ enforces a larger margin on minority-class samples and has many of its own theoretical properties Cao et al. (2019): in particular, it has been shown to counteract some of the negative effects of training with $\gamma>0$ early in training Kini et al. (2021). Although there are intuitive and theoretical rationale for each of these hyperparameters Kini et al. (2021), it is not clear how sensitive a model’s performance is to their values, especially in terms of ROC curves on a binary dataset.
@@ 4.2 Correlation between hyperparameter values and scalar metrics
To test this, we fit polynomials between the three hyperparameters ($\Omega,\gamma,\tau$) and two scalar metrics: AUC and overall accuracy on a balanced test set when $t=0.5$. Although TPR, equivalently recall, is not a good indicator of a model’s overall performance by itself (because it is trivial to increase TPR at the expense of FPR), we also fit a polynomial to TPR to gain insights on the effect of the hyperparameters on an ROC curve. We show results, including the $R^{2}$ of these regressions, in the three plots in Figure 3. We also fit three (one for AUC, accuracy, and TPR) degree-2 polynomials using all three hyperparameters as features and report $R^{2}$ values of these polynomials in the table in Figure 3. We choose polynomials of degree two instead of linear regression because we expect that an optimum exists.
We see that no single hyperparameter is strongly correlated with AUC; however, 11% of the variance can be explained by the polynomial fit to all three hyperparameters. This suggests that some variance can be removed by choosing strong hyperparameters; however, much of the variance comes from other sources. Additionally, only 60% of the variance in overall accuracy can be explained by the polynomial with all three hyperparameters. While this is significantly better than AUC’s $R^{2}$, it is still far from perfect. Thus, VS Loss hyperparameter values do not affect performance metrics in a very predictable way. It would be preferable to reduce the variance of VS Loss over performing extensive hyperparameter tuning.
@@ 4.3 Connection between hyperparameter values and softmax outputs
The third plot in Figure 3 shows that $\tau$ is strongly correlated with TPR when the threshold $t=0.5$ ($R^{2}=0.95$). To understand this, recall that $\tau$ affects the additive term on the logits in Equations 2 and 3. Specifically, increasing $\tau$ enforces a larger margin on the minority class. In general, this will lead to increasing the softmax scores of the minority class at inference time, which is similar to post-hoc VS calibration Zhao et al. (2020). Thus, when $t$ is constant, increasing $\tau$ will likely cause an increase in TPR.
We can also visualize the effect of $\tau$ by considering the set of “break-even” logits, which are defined as the set of points of $\mathbf{z}$ where the loss is equal whether the sample has label 0 or 1 (i.e., the points of $\mathbf{z}$ such that $\ell_{VS}(1,\mathbf{z})=\ell_{VS}(0,\mathbf{z})$). With regular Cross-entropy, this set is $z_{1}=z_{0}$. Assume instead that we keep $\Omega=0.5,\gamma=0$, but have arbitrary values for $\tau$, then the equation for the set of break-even points becomes $z_{1}=z_{0}+\tau\log\beta$. In other words, varying $\tau$ is equivalent to shifting the set of break-even points (and the rest of the loss landscape) by $\tau\log\beta$ [^3] Varying $\tau$ also shifts the loss landscape in a similar way for any fixed $\Omega\in(0,1)$ and $\gamma>0$. See Appendix G for details.. Figure 4 visualizes this.
In addition, $\Omega$ and $\gamma$ also have well-understood effects on the logit outputs. Appendices G and H include a general equation for the set of break-even points and more loss contours. Just as we saw in the example with $\tau$, different combinations of hyperparameters have different effects on how $\mathbf{z}$ is scaled, which in turn affects their softmax scores. Thus, training with different values of VS Loss hyperparameters corresponds to optimizing for different tradeoffs of TPR and FPR.
Figure 4: Effect of $\tau$ on loss landscape. Each plot shows $\ell_{VS}(1,\mathbf{z})-\ell_{VS}(0,\mathbf{z})$ over $z_{0},z_{1}\in[-5,5]$ for $\beta=10$. White: “Break-even” points. $\bm{\tau}$ shifts the loss landscape.